# Safe and Responsible Large Language Model : Can We Balance Bias Reduction and Language Understanding in Large Language Models?

### v3

@@ 6 Conclusion
CRediT authorship contribution statement
@@
Oluwanifemi Bamgbos: Investigation, Conceptualization, Methodology.
@@
Syed Raza Bashir: Validation, Experiments, Writing – review & editing.Data availability
Data is made available with this work.
Acknowledgements
Resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute.
@@ Appendix A Annotation Guidelines
In the development of this guide, a dedicated group of 15 annotators volunteered their expertise and time to ensure the highest standards of accuracy and sensitivity in identifying unsafe content generation. This diverse team consisted of five experts in fields related to computer science, language, psychology, and ethical computing, each accompanied by three students. An annotation guide is designed for the annotators for identifying and effectively neutralizing instances of toxicity, stereotyping, bias, and harm in textual content. This guidelines covers a diverse array of target groups and individuals, ranging from those in different age groups – children, teenagers, adults, and seniors – to individuals with varying educational backgrounds, geographic locations, and occupations, including healthcare professionals, engineers, teachers, and artists. Our aim with this annotation guideline is to ensure that all individuals, regardless of their background or characteristics, can engage with content that adheres to the principles of fairness, inclusivity, and respect.
Below is a list of target groups for which unsafe language generation happens that need to consider when annotating textual content and creating benign alternatives.
@@
Miscellaneous: Ageism, Disabled, Diseased, Feminist, Physical Appearance, Socio-economic Status.Following are the examples for annotation guidelines.
@@
Table 9: Summary of Fleiss’ Kappa Scores and Interpretations Category Fleiss Kappa Score Interpretation (Agreement) Toxicity 0.75 Substantial Harm 0.72 Substantial Bias 0.78 Substantial (-ve) Sentiment 0.70 Substantial Benign variation 0.78 Substantial The Fleiss’ Kappa scores, a measure of inter-annotator agreement, were thoroughly calculated for each category. For toxicity, we observed a Fleiss’ Kappa score of 0.75, indicating strong agreement. The harm category showed a substantial consensus with a score of 0.72. Bias identification and removal also achieved a high degree of agreement, with a score of 0.78. Negative sentiment analysis resulted in a good agreement score of 0.70. The benign generation score is calculated to be 0.78.
@@ B.1 Data Schema
Listing 1: Dataset Schema {"$schema": "http://json-schema.org/draft-07/schema#","title": "Dataset","description": "A␣dataset␣schema␣for␣analyzing␣sentences␣for␣bias,␣toxicity,␣sentiment,␣harm,␣and␣a␣benign␣variation␣for␣a␣unsafe␣text.","type": "object","properties": {"Original␣Sentence": {"type": "string","description": "The␣text␣of␣the␣original␣sentence␣that␣may␣be␣unsafe"},"BIAS": {"type": "string","description": "Indicates␣if␣the␣sentence␣is␣biased.","enum": ["Yes", "No"]},"TOXICITY": {"type": "string","description": "Level␣of␣toxicity␣of␣the␣sentence.","enum": ["No", "Mild", "High"]},"SENTIMENT": {"type": "string","description": "Sentiment␣of␣the␣sentence.","enum": ["Negative", "Neutral", "Positive"]},"HARM": {"type": "string","description": "Level␣of␣harm␣of␣the␣sentence.","enum": ["Low", "Medium", "High"]},"DEMOGRAPHIC␣TARGETING": {"type": "string","description": "Indicates␣if␣specific␣demographics␣are␣targeted.","enum": ["None", "Specific␣Demographics"]},"WORDS␣OR␣PHRASES": {"type": "string","description": "List␣of␣biased␣words␣or␣phrases␣identified␣in␣the␣sentence."},"Benign": {"type": "string","description": "The␣benign␣version␣of␣the␣text."},"required": ["Original␣Sentence", "BIAS", "TOXICITY", "SENTIMENT", "HARM", "Benign"]}B.2 Data Analysis
Table 10: Descriptive Statistics for Text Length Statistic char_length word_length Count 16614.000 16614.000 Mean 370.336 68.277 Std 451.732 80.136 Min 7.000 1.000 25% 110.000 22.000 50% 238.000 44.000 75% 491.000 90.000 Max 5000.000 1108.000 Table 10 shows the text length statistics indicate moderate average lengths but with considerable variability, suggesting the dataset contains a diverse array of text sizes, which is advantageous for creating versatile NLP models. The labels distribution is in Appendix11.
Label Category (Count) BIAS No (14227), Yes (5772) TOXICITY No (12040), Mild (5293), High (2666) SENTIMENT Negative (9028), Neutral (8370), Positive (2601) HARM Low (14151), Medium (3932), High (1915) Table 11: Summary of Label Distributions
Figure 6: Flesch Reading Ease Scores and its Relationship with Sentiment Label Figure 6 Flesch Reading Ease scores imply that positive sentiments are found in more straightforward texts, while negative and neutral sentiments appear in more complex ones, potentially affecting the ease of automated readability assessment.
Statistic Value Count 8307.000000 Mean 0.087286 Std 0.148290 Min 0.000000 25% 0.028708 50% 0.052493 75% 0.084249 Max 1.000000 Table 12: Analyzing the Impact of Benign Table 12 demonstrates the benign analysis of text alterations post-de-biasing, revealing a notable variation in the extent of change. This highlights the complexity of the debiasing process and the need for careful consideration in maintaining the original text content.
@@ Appendix D Carbon Footprint Estimation
The process for calculating energy consumption in kWh encompasses the following steps: - Total Power Consumption: Summation of the GPUs’ and CPUs’ power usage in Watts.
@@
- Calculated Energy Use (kWh): Multiplication of the total power usage (in kW) by the session duration in hours, yielding the total energy consumption.Carbon Emission Estimation:

### v5

@@ 6.3 Future Directions
While the proposed method has shown effectiveness in enhancing the model’s safety and knowledge retention, it might be worthwhile to assess its impact on the model’s general capabilities in various NLP tasks. Future work should focus on conducting comparative evaluations using datasets like Massive Multitask Language Understanding (MMLU) [53]. These evaluations will provide insights into whether the safety improvements lead to any significant trade-offs in the model’s general performance.
In future work, we aim to enhance the reliability of our bias assessment by conducting a larger number of trials and employing advanced statistical techniques, such as chi-square tests, to robustly analyze gender representation in model outputs. This will enable a more definitive conclusion about the presence of gender or other biases.
@@ 7 Conclusion
Acknowledgements Resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute.
Declarations
Funding: No funding is available.
Conflict of Interest: The authors have no conflict of interest to declare that are relevant to the content of this article.
Ethics Approval and Consent to Participate: Not applicable.
Consent for Publication: Not applicable.
Data Availability: Data is made publicly available in Huggingface.
Code Availability: The code is made publicly available at GitHub.
Author Contribution:
@@
Oluwanifemi Bamgbos: Investigation, Methodology.
@@
Syed Raza Bashir: Validation, Experiments, Writing – review & editing.
@@ Appendix A Annotation Guidelines
In the development of this guide, a dedicated group of 20 annotators volunteered their expertise and time to ensure the highest standards of accuracy and sensitivity in identifying unsafe content generation. This diverse team consisted of five experts in fields related to computer science, language, psychology, and ethical computing, each accompanied by four students. An annotation guide is designed for the annotators for identifying and effectively neutralizing instances of toxicity, stereotyping, bias, and harm in textual content. This guidelines covers a diverse array of target groups and individuals, ranging from those in different age groups – children, teenagers, adults, and seniors – to individuals with varying educational backgrounds, geographic locations, and occupations, including healthcare professionals, engineers, teachers, and artists. Our aim with this annotation guideline is to ensure that all individuals, regardless of their background or characteristics, can engage with content that adheres to the principles of fairness, inclusivity, and respect. Below is a list of target groups for which unsafe language generation happens that need to consider when annotating textual content and creating benign alternatives.
@@
Miscellaneous: Ageism, Disabled, Diseased, Feminist, Physical Appearance, Socio-economic Status.
Following are the examples for annotation guidelines.
@@
| Category | Fleiss Kappa Score | Interpretation (Agreement) |
| Toxicity | 0.75 | Substantial |
| Harm | 0.68 | Substantial |
| Bias | 0.62 | Substantial |
| (-ve) Sentiment | 0.70 | Substantial |
| Benign variation | 0.78 | Substantial |
The Fleiss Kappa scores, a measure of inter-annotator agreement, were thoroughly calculated for each category. For toxicity, we observed a Fleiss’ Kappa score of 0.75, indicating strong agreement. The harm category showed a substantial consensus with a score of 0.72. Bias identification and removal also achieved a high degree of agreement, with a score of 0.78. Negative sentiment analysis resulted in a good agreement score of 0.70. The benign generation score is calculated to be 0.78.
@@ B.1 Data Schema
B.2 Data Analysis
| Statistic | char_length | word_length |
| Count | 16614.000 | 16614.000 |
| Mean | 370.336 | 68.277 |
| Std | 451.732 | 80.136 |
| Min | 7.000 | 1.000 |
| 25% | 110.000 | 22.000 |
| 50% | 238.000 | 44.000 |
| 75% | 491.000 | 90.000 |
| Max | 5000.000 | 1108.000 |
Table 11 shows the text length statistics indicate moderate average lengths but with considerable variability, suggesting the dataset contains a diverse array of text sizes, which is advantageous for creating versatile NLP models. The labels distribution is in Appendix12.
| Label | Category (Count) |
| BIAS | No (14227), Yes (5772) |
| TOXICITY | No (12040), Mild (5293), High (2666) |
| SENTIMENT | Negative (9028), Neutral (8370), Positive (2601) |
| HARM | Low (14151), Medium (3932), High (1915) |
Figure 6 Flesch Reading Ease scores imply that positive sentiments are found in more straightforward texts, while negative and neutral sentiments appear in more complex ones, potentially affecting the ease of automated readability assessment.
| Statistic | Value |
| Count | 8307.000000 |
| Mean | 0.087286 |
| Std | 0.148290 |
| Min | 0.000000 |
| 25% | 0.028708 |
| 50% | 0.052493 |
| 75% | 0.084249 |
| Max | 1.000000 |
Table 13 demonstrates the benign analysis of text alterations post-de-biasing, revealing a notable variation in the extent of change. This highlights the complexity of the debiasing process and the need for careful consideration in maintaining the original text content.
@@ Appendix D Carbon Footprint Estimation
The process for calculating energy consumption in kWh encompasses the following steps:
- Total Power Consumption: Summation of the GPUs’ and CPUs’ power usage in Watts.
@@
- Calculated Energy Use (kWh): Multiplication of the total power usage (in kW) by the session duration in hours, yielding the total energy consumption.
Carbon Emission Estimation: